OpenStudy (anonymous):

The position of a particle in motion in the plane at time t is X(t) = −7 t I + sin(−3 t) J. At time any t, determine the following: the unit tangent vector to X(t) is: ??? I + ??? J (Note that, a unit vector is a vector whose length is 1. Multiplying any non-zero vector V by 1/|V| produces a unit vector, and multiplying any unit vector by -1 shows that there are two unit vector which are multiples of any non-zero vector. The unit tangent vector to X(t) is defined as V(t)/|V(t)|. )

OpenStudy (anonymous):

r'(t)=-7i-3cos(-3t) |r'(t)|=\[\sqrt{49+9\cos^2(-3t)}\] -7/\sqrt{49+9\cos^2(-3t)}i-3\cos(-3t)/\sqrt{49+9cos^2(-3t)k

OpenStudy (anonymous):

-7/sqrt{49+9cos^2(-3t)}i-3cos(-3t)/sqrt{49+9cos^2(-3t)k